OpenStudy (anonymous):

a tank can be filled using pipes A,B or both. It takes pipe A, running alone 18 hours to feel the tank. It takes both pipes running together 9.9 hours to fill the tank. how long does it take pipe B, running alone, to fill the tank?

OpenStudy (mertsj):

\[\frac{1}{18}+\frac{1}{x}=\frac{1}{9.9}\]

OpenStudy (anonymous):

so now di I find a common denominator?

OpenStudy (mertsj):

Multiply the 3 denominators together.

OpenStudy (anonymous):

but they don't equal the same I got 54 and 27.7

OpenStudy (anonymous):

29.7 im really confused

OpenStudy (mertsj):

\[\frac{1}{18}\times 18(9.9)(x)+\frac{1}{x}\times 18(9.9)(x)=\frac{1}{9.9}\times18(9.9)(x)\] \[9.9x+18(9.9)=18x\] \[178.2=8.1x\]

OpenStudy (mertsj):

You can finish it.

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

Alternatively, calculate the two fractions, and take their difference, which will give you 1/x, then use x = 1/(1/x)